OpenStudy (anonymous):

solve for x 1. x^5=31 2. x^4/3=-64 3. 27^2/3=x^1/2

OpenStudy (anonymous):

calculator for this one

OpenStudy (anonymous):

i'll take 3.

OpenStudy (anonymous):

unless you mean \[x^5=32\] in which case \[x=2\] because \[2^5=32\]

OpenStudy (anonymous):

27^(2/3) is the same as 3^2 is the same as 9 .. so 9 is left side of equation

OpenStudy (anonymous):

so 9=x^(1/2) square both sides to find x x=81 ..............ans.3.

OpenStudy (anonymous):

nope, it's 31 aha!

OpenStudy (anonymous):

for the first one: x=31^(1/5).............ans1

OpenStudy (anonymous):

and for the second one first take the cube root of -64=-4 then raise it to the power of 4 which is 256 x=256

OpenStudy (anonymous):

Thanks bro :)

OpenStudy (anonymous):

wait i think i did the second one wrong! however, there is something wrong with the question

OpenStudy (anonymous):

well, that's what the teacher has on the paper, -__-"

OpenStudy (anonymous):

you have to first take the 4th root of -64 and you can't take an even root of a negative number

OpenStudy (anonymous):

unless you are dealing with complex numbers.

OpenStudy (anonymous):

yes, not likely something raised to the fourth power will be negative